. How many grams of (g) can be stored in a 10.0 L container at 1000 kPa and 30C?



1
Expert's answer
2022-03-24T19:01:33-0400

Solution

Data Given;

Convert degrees to K

"Temperature = 30\u00b0C = 30 + 273 = 303K"

"Pressure = 100Kpa = 100\u00d71000 = 100000pa"


Conversion of Pascal pressure to Atmospheric pressure

Let the unknown pressure be = x

"1 atm = 1.01325\u00d710^5pa"


x atm = 100000pa

"x = 100000\/1.01325\u00d710^5 = 0.9869 atm"

Pressure = 0.9869 atm

Therefore,

100Kpa = 100000pa = 0.9869 atm


From general gas equation,

PV = nRT

Where P = Pressure

V = Volume

n = Number of moles

"R = Constant = 0.0821 atm L\/ mole K"

T = Temperature


PV = nRT

n = PV/RT

"n = 0.9869×\frac{10}{0.0821}×303"


n = 9.869/24.8763 = 0.3967 moles


Recall that,

Number of moles (n) Mass/Molar mass

Mass = n × Molar mass

But Molar mass of Cl2 "= 35.5\u00d72" = 71 g/mole

"Mass = 0.3967\times71"

= 28.1657 ≈ 28.17g


The mass of Chlorine that can be stored in the container is = 28.17g